OpenStudy (anonymous):

1. Given a triangle with leg A = 12 and the hypotenuse C = 20, find the the length of leg B.

OpenStudy (michele_laino):

please, is your triangle a right triangle?

OpenStudy (anonymous):

that is the full question

OpenStudy (michele_laino):

Since I see the term "hypotenuse", I think that it is a right triangle

OpenStudy (michele_laino):

So in order to find B, you can apply the Pitagora theorem, namely: \[B ^{2}=C ^{2}-A ^{2}\]

OpenStudy (anonymous):

can you please solve it so i can see how you do it

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

\[B ^{2}=20^{2}-12^{2}=400-144=256\] so, we have: \[B=16\] since 16^2=256

OpenStudy (michele_laino):



OpenStudy (michele_laino):